OpenStudy (anonymous):

Algebra 2 help, please?

OpenStudy (anonymous):

\[If f(x)=3x+1 and f ^{-1}=\frac{ x-1 }{ 3 } then f ^{-1}(7)=\]

OpenStudy (dtan5457):

You would just plug 7 as the x value of (f)^-1

OpenStudy (dtan5457):

You already have the inverse of f(x), they want f(^-1)(7)

OpenStudy (anonymous):

A.) 22 B.) 2 C.) -2

OpenStudy (dtan5457):

Did you try plugging 7 into the inverse function? It becomes straight forward once you do

OpenStudy (anonymous):

I don't know how to do these equations.

OpenStudy (anonymous):

I still don't get it..lol

OpenStudy (dtan5457):

I had a typo it's \[f^{-1}(7)=\frac{ (7)-1 }{ 3 }, or \frac{ 6 }{ 3 }\]

OpenStudy (dtan5457):

6/3 is the most simplified way i can possibly explain it to you

OpenStudy (dtan5457):

6/3=?

OpenStudy (anonymous):

\[if f(x)=3x+1 and f ^{-1}=\frac{ x-1 }{ 3 } then the ordered pair of ƒ(3) = (3, 10) (3, 2/3) (10, 3)\]

OpenStudy (dtan5457):

Can you screen shot? your question is cut off., don't write words into equation form lol

OpenStudy (anonymous):

I'm on my laptop..lol

OpenStudy (dtan5457):

i can't understand your question lol laptops can take screenshots too